{2x + y=7
{-5x - 2y=8

Answer by RAY100(1637)

1) 2x +y = 7
2) -5x -2y =8
.
(1) *2 = 4x +2y =14,,,,,(1a)
.
add (1a) + (2) = -x = 22,,,,,x=-22
.
subst in (1)
.
2(-22) + y = 7
.
y = 51
.
checking
(1) 2(-22) + 51 = 7,,,ok
(2) -5(-22) - 2(51) = 110-102=8,,,,ok